Please help me solve this system using the method of substitution. I would like the steps to know how it's done. Thank you.
.5x+2.2y=9
6x+.4y=22

Answer by stanbon(75887)

5x+2.2y=9
6x+.4y=22
----
Solve the 1st for "x":
x = (9-2.2y)/5
----
Substitute for "x" and solve for "y":
6[(9-2.2y)/5] + 0.4y = 22
------------------------------
(54 - 13.2y)/5 + 0.4y = 22
Multiply thru by 5 to get:
54-13.2y + 2y = 110
-11.2y = 56
y = -5
---
Solve for "x":
x = (9-2.2y)/5
x = (9-2.2*-5)/5
x = 4
==============
Cheers,
Stan H.
================